BlockOut Co. has 71,362 bonds outstanding that are selling at par value. The bonds yield 8.5 percent. The company also has 4.2 million shares of common stock outstanding. The stock has a beta of 1.19 and sells for $40.4 a share. The U.S. Treasury bill is yielding 4.3 percent and the market risk premium is 8.5 percent. Blackout's tax rate is 32 percent. What is the firm's weighted average cost of capital? (Enter answer in percents.)
Answer #1

Cost of Debt = 8.50%

Value of Debt = $71,362,000

Value of Equity = 4,200,000(40.40) = $169,680,000

As per CAPM Model,

Cost of Equity = 0.043 + 1.19(0.085)

Cost of Equity = 14.42%

Weight of Debt = 71,362,000/(71,362,000 + 169,680,000)

Weight of Debt = 0.2961

Weight of Equity = 1 - 0.2961 = 0.7039

WACC = 0.2961(1 - 0.32)(0.085) + 0.7039(0.1442)

WACC = 11.86%